# Decentralized Reliability Estimation for Mixnets

### v1

@@ 3.2 VRF-Based Routing
Proof of no-skipping. For any counter value $ctr$ that does not correspond to a measurement packet, a gateway can open the corresponding VRF value that determines this particular opportunity is a non-measurement legitimate packet. The gateway reveals the tuple $\langle ctr,r_{\mathsf{type}},\pi_{\mathsf{type}}\rangle$ which can be verified against $vk_{\mathsf{vrf}}$ in conjunction to the test $r_{\mathsf{type}}\geq T$. In a setting where all measurement packets have been opened, a random subset of the complement set of non-measurement counters can be challenged in this way to ensure that no measurement packets have been skipped. In more detail, if $N$ is the number of proposed non-measurement packets, by opening $v$ VRF values, the probability of missing all $v$ times a skipped measurement packet is less than $(1-1/N)^{v}<\exp(-v/N)$. It follows that the probability of finding it is at least $1-\exp(-v/N)=\alpha_{\mathsf{ns}}$, i.e., we need $v=-N\ln(1-\alpha_{\mathsf{ns}})$. In terms of size of a single opening, observe that we need to reveal one index, one group element one hash and two exponents for the proof. Using Ed25519, this results to 132 Bytes per counter position.
@@
The first key observation is that conditioning on a legitimate packet being sent by the entry gateway, we obtain a coin flip with probability $p_{\mathsf{lot}}$ for sending a measurement packet. Due to gateways being utility maximizers observe that the total number of announced packets sent by gateway $g$, $N^{\mathsf{gw}}_{g}=\sum_{c}N_{g,c}$ would be at least as large as the actual number of packets sent.
@@
We next require that entry gateways open all the measurement packets in the announced ranges for each credential. Due to the proof of no skipping, any gateway that does not want to get caught lying must open all the measurement packets in the range for each credential it announced. Moreover, note that if there is a credential that is misreported to have been used less than in reality, e.g., the declared $N_{g,c}$ is smaller than the actual $N^{\mathsf{a}}_{g,c}$, then there should be another credential $c^{\prime}$ that compensates for it by the same gateway, i.e., it holds $N^{\mathsf{a}}_{g,c^{\prime}}<N_{g,c^{\prime}}$. It follows that for credential $c^{\prime}$ there will be measurement packets that have not being opened and hence this exposes the gateway $g$ who will be caught in the proof of no skipping for credential $c^{\prime}$.
@@
Furthermore: (1) If measurement packets are not sent (despite being opened) this will reduce the performance of the gateway, so a utility maximizing gateway will not perform this action. (2) if non-measurement legitimate packets are selectively not sent (despite having the opportunity to send them) there are two possibilities: (i) the non-measurement packets are simply dropped, in which case client satisfaction drops and thus the gateway will not prefer this strategy; (ii) a gateway that colludes with the client may substitute legitimate packets by illegitimate packets that carry the same payload — this does not affect our estimation measurements that are only concerned with performance on legitimate packets.
@@
For receiving gateways, assuming they are utility maximizers, we observe: (1) if measurement packet tags are not recorded by the gateway in the Bloom filter despite being received, the performance will drop, hence the gateway will not perform this action; (2) if non-measurement packets are not handed over to their destination this means that they have been processed, detected as non-measurement and then dropped. This affects client satisfaction (e.g, consider the case of clients sending self-loops) and as a result is not a preferred action of a utility maximizer gateway.
4 Node performance estimation
@@
The goal of this protocol is to estimate, for every intermediary node $j$ in the mixnet, a performance score $\hat{\rho}_{j}$ that reflects the rate of packet losses caused by $j$, with $\hat{\rho}_{j}=0$ meaning that $j$ lost all packets sent to it (e.g., if $j$ is offline the entire epoch) and $\hat{\rho}_{j}=1$ that the node was consistently online and correctly processed and sent all received packets. These $\hat{\rho}_{j}$ can be computed by any entity that knows the outputs of the sampling protocol described in Sect. 3, i.e., the amounts ${s}^{*}_{e}$ and ${d}^{*}_{e}$ of transmitted and dropped measurement packets per link $e\in G$. We consider reliable link transmissions such that any packet drops are either due to a failure of the predecessor or the successor. The estimated node performance $\hat{\rho}_{j}$ is computed as:
$\hat{\rho}_{j}=\frac{\sum_{k\in S(j)}{s}^{*}_{(j,k)}+\hat{\beta}_{(j,k)}{d}^{*}_{(j,k)}}{\sum_{i\in P(j)}{s}^{*}_{(i,j)}+\hat{\beta}_{(i,j)}{d}^{*}_{(i,j)}}$ (7)
@@
In the case of gateways, $\hat{\rho}_{g}$ is computed considering the successfully transmitted (and dropped) measurements to first-layer successors $i\in S(g)$ ($g$’s outgoing links); and the received (and dropped) measurements from last-layer predecessors $k\in P(g)$ recorded in their Bloom filter ($g$’s incoming links).
@@
The parameter $\hat{\beta}_{e}$ denotes the fraction of drops attributed to the link successor, with the remaining $1-\hat{\beta}_{e}$ being attributed to the link predecessor. Ideally, participants verify that $\hat{\beta}_{e}\approx{\beta}_{e}$, where ${\beta}_{e}$ is the actual fraction o drops caused by the link successor. In a network deployment where packets are transmitted via point-to-point links $e=(i,j)$, it is however impossible for third parties to ascertain the true value of ${\beta}_{e}$, and thus $\hat{\beta}_{e}$ must be estimated based on knowledge on the patterns of node failure and the available evidence in the entire network. We thus consider the observable behavior of the two counterparties across all of their connections adopting a threshold approach to estimate $\hat{\beta}_{e}$.
@@
Given the link performance estimates $\hat{\rho}_{e}$ for $e\in G$, computed according to Eq. 4, for a node $j$ we define $\bar{\rho}^{\mathsf{in}}_{j}$ as the median value of the incoming set of $\hat{\rho}_{(i,j)}$ for $i\in P(j)$, while $\bar{\rho}^{\mathsf{out}}_{j}$ is the median value of the outgoing set of $\hat{\rho}_{(j,k)}$ for $k\in S(j)$. These median values for the incoming and outgoing edge sets are compared to a preset threshold $\bar{\tau}$. If node $j$’s median incoming (resp. outgoing) performance $\bar{\rho}^{\mathsf{in}}_{j}\geq\bar{\tau}$ (resp. $\bar{\rho}^{\mathsf{out}}_{j}\geq\bar{\tau}$), then $j$ is considered to be functioning reliably on its input (resp. output) links; and unreliably otherwise. In this way, each node $j$ receives a pair of labels, one for the set of input edges and one for the output set, each independently set to either reliable or unreliable depending on the node’s median edge performance in each direction.
@@
Given a link $e=(i,j)$, if both $i$’s output and $j$’s input are reliable, or both are unreliable, then any drops ${d}^{*}_{(i,j)}$ are symmetrically attributed to $i$ and $j$ (half to each with $\hat{\beta}_{(i,j)}=\frac{1}{2}$). If on the other hand one of the two is reliable while the other is unreliable, then the one identified as unreliable is attributed all the drops in that edge ($\hat{\beta}_{(i,j)}=0$ if $i$’s output is unreliable and $\hat{\beta}_{(i,j)}=1$ if $j$’s input is unreliable). This mechanism effectively identifies unreliable nodes subject to failures such as going offline or lacking enough throughput, as these failures affect the node’s median performance. Additionally, in adversarial settings the malicious nodes engaging in selective packet dropping suffer as much performance degradation as they are able to inflict on their (otherwise reliable) targets.
@@
We refer to Appendix C for further details on the rationale and implications of various approaches to assigning values to $\hat{\beta}_{e}$, and additional explanations why the threshold approach outlined above (and in more detail in Appendix C.3) offers the best tradeoffs.
@@
Performance estimation in unreliable settings. We first consider scenarios where all nodes are non-adversarial but up to half of each layer may be unreliable, go offline, have limited throughput or otherwise be subject to random failures, and argue that the estimated $\hat{\rho}_{j}$ is a good approximation of the underlying ${\rho}_{j}$ (given by Eq. 7 but using ‘ground truth’ ${s}_{e}$, ${d}_{e}$ and ${\beta}_{e}$ instead of ${s}^{*}_{e}$, ${d}^{*}_{e}$ and $\hat{\beta}_{e}$).
@@
In case $j$ is unreliable, observe that for reliable predecessors $i$ and successors $k$, it also holds ${\beta}_{(i,j)}=1=\hat{\beta}_{(i,j)}$ and ${\beta}_{(j,k)}=0=\hat{\beta}_{(j,k)}$, which accurately attributes packet drops to $j$. On the other hand, if either $i$ or $k$ are unreliable as well, a disparity between downtimes with $j$ can lead to a small deviation in the estimation $\hat{\rho}_{j}$ with respect to ${\rho}_{j}$ (since the algorithm will set the $\hat{\beta}_{(i,j)}$ or $\hat{\beta}_{(j,k)}$ to $1/2$ thus averaging out both sub-par performances). We investigate this experimentally in Sect. 4.1.1 and demonstrate that we still obtain a good approximation $\hat{\rho}_{j}\approx{\rho}_{j}$ of the node performance score, with an accuracy that increases with the number of measurement packets.
Performance estimation in adversarial settings. We next discuss an adversarial setting where a colluding set of malicious nodes may drop packets to and from honest nodes that are the targets of the attack. The purpose of the attack is to cause the estimated performance of a target $j$ (or set of targets) to be lower than its actual performance, i.e., $\hat{\rho}_{j}<{\rho}_{j}$.
@@
Rather than obtaining an immediate benefit, an adversary engaging in such an attack may ultimately seek to harm the target’s reputation to create opportunities for his own nodes [13]. For example, in Nym both the estimated performance and the amount of stake delegated to a node determine node participation and rewards. A lower estimated performance negatively affects node profitability, driving drive away the node’s stake delegations – which adversaries may hope to attract for their own nodes. We aim to cause an adversary deploying this attack an aggregate degradation in estimated performance of adversarial nodes at least on par to the aggregate degradation suffered by the targets. If this is the case, the adversary fails to achieve its ultimate goal, since adversarial nodes will also exhibit poor performance and will also lose, rather than increase, their rewards, reputation and attractiveness to staking delegates, paying a price for engaging in the attack without any short-term or long-term benefits.
@@
Consider $j$ an honest reliable node being the target of the attack. Observe that as long as malicious nodes are less than half of a layer preceeding or succeeding $j$, the best the adversary can do is drop packets directed to $j$ in a way that both counterparties in the link will get the blame. This is due to the fact that the adversary cannot make $j$ appear unreliable (since $j$ is connected to a majority of other reliable nodes and hence its median values $\bar{\rho}^{\mathsf{in}}_{j},\bar{\rho}^{\mathsf{out}}_{j}$ are above the threshold). It follows that the best strategy for the adversary is to ensure their nodes are reliable (by maintaining good performance in a majority of incoming and outgoing links) and lead the algorithm to the equitable $\hat{\beta}=1/2$ choice of sharing blame between the two counterparties in each link it shares with $j$. In Sect. 4.1.2 we demonstrate experimentally this intuition and the proportionality of performance degradation between the adversary and the target nodes.
@@ 4.1 Simulation-based empirical evaluation
We have implemented a discrete-event mixnet simulator to empirically evaluate whether the proposed approach provides accurate performance measurement estimations in different scenarios. Simulation allows us to model various network conditions including reliable, unreliable and adversarial node behaviors, for which we know the ground truth, and compare the nodes’ true performance ${\rho}_{j}$ to the performance $\hat{\rho}_{j}$ that is estimated based on the available measurement samples. We conduct an empirical evaluation of the estimation error ${\epsilon}_{j}=\hat{\rho}_{j}-{\rho}_{j}$ per node $j$, in both unreliable and adversarial settings. Appendix D describes in detail the implemented simulator, its outputs, the experimental setups used for the evaluation in unreliable and adversarial settings, and additional results for scenarios that combine both unreliable and adversarial participants.
@@ 4.1.1 Unreliable setting
We first evaluate the proposed node performance estimation protocol in scenarios where all nodes are honest but up to half the nodes in every layer are unreliable, as described in Appendix D.5. To evaluate the accuracy of the estimation given different amounts of measurement samples, we consider seven setups with between $25$k and $2$ million measurement packets in the epoch, corresponding to between $2.5$ and $200$ million total encoded packets for $p_{\mathsf{lot}}=1\%$. In each simulation run we compute ${\epsilon}_{j}=\hat{\rho}_{j}-{\rho}_{j}$ for every routing intermediary $j$. Figure 2 shows the resulting distributions of ${\epsilon}_{j}$ in each setup (aggregating samples from $20$ runs), distinguishing between reliable (black) and unreliable (blue) nodes. Note that ${\rho}_{j}=1$ for reliable nodes, and thus ${\epsilon}_{j}\leq 0$ for these, since their performance cannot be overestimated. For unreliable nodes, i.e., such that ${\rho}_{j}<1$, the error ${\epsilon}_{j}$ can be positive or negative.
@@
As we can see in the figure, the performance estimation error is small for reliable nodes (black boxes), and it becomes negligible once there are $100$k or more measurement samples. In such scenarios, the proposed node performance estimation protocols are thus remarkably accurate in their estimations for reliable nodes. The estimation error is larger for unreliable nodes, though it decreases as more measurement samples are available: with $25$k samples the error range exceeds $7\%$ of possible performance over-estimation and $5\%$ of under-estimation, while this error range is within $1\%$ of under- or over-estimation when $2$ million measurements are used. Note that nodes are rearranged in the mixnet per epoch and measured independently in each epoch. Thus, we can expect that per-epoch performance estimation errors cancel each other out when averaging node performance over multiple epochs.

### v2

@@ 3.2 VRF-Based Routing
Proof of no-skipping. For any counter value $ctr$ that does not correspond to a measurement packet, a gateway can open the corresponding VRF value that determines this particular opportunity and demonstrate it corresponds to a non-measurement legitimate packet. Namely, the gateway reveals the tuple $\langle ctr,r_{\mathsf{type}},\pi_{\mathsf{type}}\rangle$ which can be verified against $vk_{\mathsf{vrf}}$ in conjunction to the test $r_{\mathsf{type}}\geq T$. In a setting where all measurement packets have been opened, a random subset of the complement set of non-measurement counters can be challenged in this way to ensure that no measurement packets have been skipped. In more detail, if $N$ is the number of proposed non-measurement packets, by opening $v$ VRF values, the probability of missing all $v$ times a skipped measurement packet is less than $(1-1/N)^{v}<\exp(-v/N)$. It follows that the probability of finding it is at least $1-\exp(-v/N)=\alpha_{\mathsf{ns}}$, i.e., we need $v=-N\ln(1-\alpha_{\mathsf{ns}})$. In terms of size of a single opening, observe that we need to reveal one index, one group element one hash and two exponents for the proof. Using Ed25519, this results to 132 Bytes per counter position.
@@
The first key observation is that conditioning on a legitimate packet being sent by the entry gateway, we obtain a coin flip with probability $p_{\mathsf{lot}}$ for sending a measurement packet. Due to gateways being utility maximizers observe that the total number of announced packets sent by gateway $g$, $S_{g}=\sum_{c}s_{(c,g)}$ would be at least as large as the actual number of packets sent.
@@
We next require that entry gateways open all the measurement packets in the announced ranges for each credential. Due to the proof of no skipping, any gateway that does not want to get caught lying must open all the measurement packets in the range for each credential it announced. Moreover, note that if there is a credential that is misreported to have been used less than in reality, e.g., the declared $s_{(c,g)}$ is smaller than the actual $s^{\mathsf{a}}_{(c,g)}$, then there should be another credential $c^{\prime}$ that compensates for it by the same gateway, i.e., it holds $s^{\mathsf{a}}_{(c^{\prime},g)}<s_{(c^{\prime},g)}$. It follows that for credential $c^{\prime}$ there are measurement packets that have not been opened and hence this exposes the gateway $g$ who will be caught in the proof of no skipping for credential $c^{\prime}$.
@@
Furthermore: (1) If measurement packets are not sent (despite being opened) this will reduce the reliability of the gateway, so a utility maximizing gateway will not perform this action. (2) If non-measurement legitimate packets are selectively not sent (despite having the opportunity to send them) there are two possibilities: (i) the non-measurement packets are simply dropped, in which case client satisfaction drops and thus the gateway will not prefer this strategy; (ii) a gateway that colludes with the client may substitute legitimate packets by illegitimate packets that carry the same payload — this does not affect our estimation measurements that are only concerned with reliability of legitimate packets.
@@
For receiving gateways, assuming they are utility maximizers, we observe: (1) If measurement packet tags are not recorded by the gateway in the Bloom filter despite being received, the reliability will drop, hence the gateway will not perform this action. (2) If non-measurement packets are not handed over to their destination this means that they have been processed, detected as non-measurement and then dropped. This affects client satisfaction (e.g, consider the case of clients sending self-loops) and as a result is not a preferred action of a utility maximizer gateway.
4 Node reliability estimation
@@
The goal of this protocol is to estimate, for every node $j$ in the mixnet, a reliability score $\hat{\rho}_{j}$ that reflects the rate of packet losses caused by $j$, with $\hat{\rho}_{j}=0$ meaning that $j$ lost all packets sent to it (e.g., if $j$ is offline the entire epoch) and $\hat{\rho}_{j}=1$ that the node was consistently online and correctly processed and sent all received packets. These $\hat{\rho}_{j}$ can be computed by any entity that knows the outputs of the sampling protocol described in Sect. 3, i.e., the amounts ${s}^{*}_{e}$ and ${d}^{*}_{e}$ of transmitted and dropped measurement packets per link $e\in G$. We consider reliable link transmissions such that any packet drops are either due to a failure of the predecessor or the successor. The estimated node reliability $\hat{\rho}_{j}$ is computed as:
$\hat{\rho}_{j}=\frac{\sum_{k\in S(j)}({s}^{*}_{(j,k)}+\hat{\beta}_{(j,k)}{d}^{*}_{(j,k)})}{\sum_{i\in P(j)}({s}^{*}_{(i,j)}+\hat{\beta}_{(i,j)}{d}^{*}_{(i,j)})}$ (7)
@@
In the case of gateways, $\hat{\rho}_{g}$ is computed considering the successfully transmitted (and dropped) measurements to first-layer node successors $i\in S(g)$ ($g$’s outgoing links); and the received (and dropped) measurements from last-layer node predecessors $k\in P(g)$ recorded in their Bloom filter ($g$’s incoming links).
@@
The parameter $\hat{\beta}_{e}$ denotes the fraction of drops attributed to the link successor, with the remaining $1-\hat{\beta}_{e}$ being attributed to the link predecessor. Ideally, participants verify that $\hat{\beta}_{e}\approx{\beta}_{e}$, where ${\beta}_{e}$ is the actual fraction of drops caused by the link successor. In a network deployment where packets are transmitted via point-to-point links $e=(i,j)$, it is however impossible for third parties to ascertain the true value of ${\beta}_{e}$, and thus $\hat{\beta}_{e}$ must be estimated based on knowledge on the patterns of node failure and the available evidence in the entire network. We thus consider the observable behavior of the two counterparties across all of their connections adopting a threshold approach to estimate $\hat{\beta}_{e}$.
@@
Given the link reliability estimates $\hat{\rho}_{e}$ for $e\in G$, computed according to Eq. 4, for a node $j$ we define $\bar{\rho}^{\mathsf{in}}_{j}$ as the median value of the incoming set of $\hat{\rho}_{(i,j)}$ for $i\in P(j)$, while $\bar{\rho}^{\mathsf{out}}_{j}$ is the median value of the outgoing set of $\hat{\rho}_{(j,k)}$ for $k\in S(j)$. These median values for the incoming and outgoing edge sets are compared to a preset threshold $\bar{\tau}$. If node $j$’s median incoming (resp. outgoing) reliability $\bar{\rho}^{\mathsf{in}}_{j}\geq\bar{\tau}$ (resp. $\bar{\rho}^{\mathsf{out}}_{j}\geq\bar{\tau}$), then $j$ is considered to be functioning reliably on its input (resp. output) links; and unreliably otherwise. In this way, each node $j$ receives a pair of labels, one for the set of input edges and one for the output set, each independently set to either reliable or unreliable depending on the node’s median link reliability in each direction.
@@
Given a link $e=(i,j)$, if both $i$’s output and $j$’s input are reliable, or both are unreliable, then any drops ${d}^{*}_{(i,j)}$ are symmetrically attributed to $i$ and $j$ (half to each with $\hat{\beta}_{(i,j)}=\frac{1}{2}$). If on the other hand one of the two is reliable while the other is unreliable, then the one identified as unreliable is attributed all the drops in that edge ($\hat{\beta}_{(i,j)}=0$ if $i$’s output is unreliable and $\hat{\beta}_{(i,j)}=1$ if $j$’s input is unreliable). This mechanism effectively identifies unreliable nodes subject to failures such as going offline or lacking enough throughput, as these failures affect the node’s median reliability. Additionally, in adversarial settings the malicious nodes engaging in selective packet dropping suffer as much reliability degradation as they are able to inflict on their (otherwise reliable) targets.
@@
We refer to Appendix A.2 for further details on the rationale and implications of various approaches to assigning values to $\hat{\beta}_{e}$, and additional explanations why the threshold approach outlined above (and in more detail in Appendix A.2.3) offers the best tradeoffs.
@@
Reliability estimation in unreliable settings. We first consider scenarios where all nodes are non-adversarial but up to half of each layer may be unreliable, go offline, have limited throughput or otherwise be subject to random failures, and argue that the estimated $\hat{\rho}_{j}$ is a good approximation of the underlying ${\rho}_{j}$ (given by Eq. 7 but using ‘ground truth’ ${s}_{e}$, ${d}_{e}$ and ${\beta}_{e}$ instead of ${s}^{*}_{e}$, ${d}^{*}_{e}$ and $\hat{\beta}_{e}$).
@@
In case $j$ is unreliable, observe that for reliable predecessors $i$ and successors $k$, it also holds ${\beta}_{(i,j)}=1=\hat{\beta}_{(i,j)}$ and ${\beta}_{(j,k)}=0=\hat{\beta}_{(j,k)}$, which accurately attributes packet drops to $j$. On the other hand, if either $i$ or $k$ are unreliable as well, a disparity between downtimes with $j$ can lead to a small deviation in the estimation $\hat{\rho}_{j}$ with respect to ${\rho}_{j}$ (since the algorithm will set the $\hat{\beta}_{(i,j)}$ or $\hat{\beta}_{(j,k)}$ to $1/2$ thus averaging out both sub-par performances). We investigate this experimentally in Sect. 4.1.1 and demonstrate that we still obtain a good approximation $\hat{\rho}_{j}\approx{\rho}_{j}$ of the node reliability score, with an accuracy that increases with the number of measurement packets.
@@
Reliability estimation in adversarial settings. Adversarial nodes that are unreliable and miss the reception of some packets addressed to them have no way to include the right evidence in their Bloom filters to pretend to have higher reliability than they actually have. An alternative attack is to cause the estimated reliability of a target $j$ (or set of targets) to be lower than its actual reliability, i.e., $\hat{\rho}_{j}<{\rho}_{j}$. We next discuss an adversarial setting where a colluding set of malicious nodes may drop packets to and from honest nodes that are the targets of the attack. Rather than obtaining an immediate benefit, an adversary engaging in such an attack may ultimately seek to harm the target’s reputation to create opportunities for his own nodes [17]. For example, in Nym the estimated reliability of a node determines its participation and rewards, and a low estimated reliability may result in the exclusion of the targeted nodes. We aim to cause an adversary deploying this attack an aggregate degradation in estimated reliability of adversarial nodes at least on par to the aggregate degradation suffered by the targets. If this is the case, the adversary fails to achieve its ultimate goal, since adversarial nodes will also exhibit poor reliability and will also lose, rather than increase, their rewards and reputation with a reliability score lower than if they did not engage in any attack, thus paying a price for engaging in the attack.
@@
Consider $j$ an honest and reliable node being the target of the attack. As long as malicious nodes are less than half of a layer preceding or succeeding $j$, the best the adversary can do is drop packets directed to $j$ in a way that both counterparties in the link will get the blame. This is due to the fact that the adversary cannot make $j$ appear unreliable (since $j$ is connected to a majority of other reliable nodes and hence its median values $\bar{\rho}^{\mathsf{in}}_{j},\bar{\rho}^{\mathsf{out}}_{j}$ are above the threshold). It follows that the best strategy for adversaries is to ensure their own nodes are reliable (by maintaining good reliability in a majority of incoming and outgoing links) and lead the algorithm to the equitable $\hat{\beta}=1/2$ choice of sharing blame between the two counter-parties in each link it shares with $j$. In Sect. 4.1.2 we demonstrate experimentally this intuition and the proportionality of reliability degradation between the adversary and the target nodes.
@@ 4.1 Simulation-based empirical evaluation
We have implemented a discrete-event mixnet simulator to empirically evaluate whether the proposed approach provides accurate reliability measurement estimations in different scenarios. Simulation allows us to model various network conditions including reliable, unreliable and adversarial node behaviors, for which we know the ground truth, and compare the nodes’ true reliability ${\rho}_{j}$ to the reliability $\hat{\rho}_{j}$ that is estimated based on the available measurement samples. Appendix A.3 describes in detail the implemented simulator, its outputs, the experimental setups used for the evaluation in unreliable and adversarial settings, and additional results for scenarios that combine both unreliable and adversarial participants.
@@ 4.1.1 Unreliable setting
We first evaluate the proposed node reliability estimation protocol in scenarios where all nodes are honest but up to half the nodes in every layer are unreliable, as described in Appendix A.3.5. To evaluate the accuracy of the estimation given different amounts of measurement samples, we consider seven setups with between $25$k and $2$ million measurement packets in the epoch, corresponding to between $2.5$ and $200$ million total encoded packets for $p_{\mathsf{lot}}=1\%$. Figure 2 shows the resulting distributions of ${\epsilon}_{j}$ in each setup (aggregating samples from $20$ runs), distinguishing between reliable (black) and unreliable (blue) nodes. Note that ${\rho}_{j}=1$ for reliable nodes, and thus ${\epsilon}_{j}\leq 0$ for these, since their reliability cannot be overestimated. For unreliable nodes, i.e., such that ${\rho}_{j}<1$, the error ${\epsilon}_{j}$ can be positive or negative.
@@
As we can see in the figure, the reliability estimation error is small for reliable nodes (black boxes), and it becomes negligible once there are $100$k or more measurement samples. In such scenarios, the proposed node reliability estimation protocols are thus remarkably accurate in their estimations for reliable nodes. The estimation error is larger for unreliable nodes, though it decreases as more measurement samples are available: with $25$k samples the error range exceeds $7\%$ of possible reliability over-estimation and $5\%$ of under-estimation, while this error range is within $1\%$ of under- or over-estimation when $2$ million measurements are used. Note that nodes are rearranged in the mixnet per epoch and measured independently in each epoch. Thus, we can expect that per-epoch reliability estimation errors cancel each other out when averaging node reliability over multiple epochs.